OpenStudy (anonymous):

Write down the vector equation for the line through the given point in the specified direction. Then eliminate t to obtain the cartesian equation. (a) (2,3), \[\left(\begin{matrix}1 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

(x,y) = (2,3) +t(1,2) x= 2+1t, y = 3+2t eliminating t: t =(x-2)/1 =( y-3)/2 2(x-2)=y-3 y = 2x -1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you wellcome